# UniFlow: Towards Zero-Shot LiDAR Scene Flow for Autonomous Vehicles via Cross-Domain Generalization

### v1

@@ Appendix D Performance Breakdown on TruckScenes
| Flow4D (UniFlow, Ours) | 23.59 | 68.41 | 1.78 | 0.57 | 0.281 | 0.088 | 0.277 | 0.530 | 0.230 | 0.034 | 0.097 | 0.119 | 0.389 |
@@
Appendix E Visualizing Failure Cases
Scene 1 |
Scene 2 |
(a) Ground truth
(b) UniFlow
(c) Camera View
Figure 6 illustrates two representative failure modes of UniFlow models. In the first example, adverse weather introduces dense rain streaks that create spurious LiDAR returns, leading the model to produce localized artifacts. In the second example, extreme sparsity at long range causes the model not estimate flow for an object located nearly 100m away.
Appendix F Supplementary Video
We include a supplementary video on our project page, where we visualize scene flow predictions on a challenging rainy sequence from TruckScenes. As shown in the video, Flow frequently produces artifacts on rain streaks and background points, which become especially pronounced during occlusions. Flow also predicts inconsistent flow vectors on dynamic objects. In contrast, Flow (UniFlow) yields significantly more stable and coherent motion fields.

### v2

@@ Appendix 0.D Ablation on Velocity Augmentation
| Flow4D-XL (UniFlow, Ours) | 23.59 | 68.41 | 1.78 | 0.57 | 0.281 | 0.088 | 0.277 | 0.530 | 0.230 | 0.034 | 0.097 | 0.119 | 0.389 |
@@
Appendix 0.E Performance Breakdown on TruckScenes
Appendix 0.F Impact of Model Size vs. Dataset Size
In Table 10, we found that scaling up Flow4D’s backbone and scaling up training data considerably improved out-of-domain performance on TruckScenes. We explicitly disentangle the impact of model size and dataset size on both in-domain and out-of-domain performance in Table 17. We find that training Flow4D on the unified dataset improves performance on all datasets except nuScenes. We note that training Flow4D-XL with the standard training recipe yields worse performance than Flow4D on AV2 and nuScenes. Intuitively, dramatically increasing model capacity without increasing training data size leads to underfitting. Lastly, we find that scaling up model size and training data achieves the best performance, improving performance across all in-domain and out-of-domain datasets.
| Methods | Dynamic Bucket-Normalized | ||||
| # Params. | AV2 | Waymo | nuScenes | TruckScenesZS | |
| Flow4D | 4.6M | 0.192 | 0.215 | 0.230 | 0.456 |
| Flow4D (UniFlow, Ours) | 4.6M | 0.189 | 0.204 | 0.257 | 0.301 |
| Flow4D-XL | 41.4M | 0.196 | 0.210 | 0.249 | 0.416 |
| Flow4D-XL (UniFlow, Ours) | 41.4M | 0.176 | 0.191 | 0.196 | 0.281 |
| TruckScenes Zero-Shot | Dynamic Bucket-Norm. | ||||
| Dyn. M | Car | Other | Pedestr. | VRU | |
| Flow (AV2) + Domain Adaptation | 0.528 | 0.366 | 0.520 | 0.555 | 0.528 |
| Flow (Waymo) + Domain Adaptation | 0.331 | 0.157 | 0.411 | 0.549 | 0.207 |
| Flow (nuScenes) + Domain Adaptation | 0.667 | 0.533 | 0.730 | 0.709 | 0.696 |
| Flow (UniFlow, Ours) | 0.283 | 0.101 | 0.293 | 0.513 | 0.226 |
Appendix 0.G Comparison with Domain Adaptation Baselines
We additionally compare UniFlow against domain adaptation baselines on TruckScenes zero-shot evaluation. Specifically, we apply the same data augmentation strategies to dataset-specific Flow models trained on AV2, Waymo, and nuScenes, and compare them against Flow trained with our unified multi-dataset pipeline. As shown in Table 18, Flow (UniFlow) achieves the best overall performance, outperforming the best domain adaptation baseline by 14.5% on Dynamic Bucket-Normalized EPE. These results suggest that unified multi-dataset training is a particularly effective strategy for improving cross-domain generalization.
Appendix 0.H Zero-Shot Visualizations on AEVAScenes
We provide additional qualitative results on AEVAScenes to illustrate zero-shot generalization to a previously unseen LiDAR modality. Figure 6 compares dataset-specific Flow, Flow (UniFlow), and ground truth on two representative scenes. In both examples, UniFlow recovers the motion of foreground vehicles that the dataset-specific baseline fails to capture, highlighted by the blue vehicles in the red circles. UniFlow also produces more accurate motion estimates on the small foreground object to the right in Scene I.
Scene I
Scene II
| Flow Flow (UniFlow) GT |
Scene I
Scene II
| GT UniFlow RGB |
Appendix 0.I Visualizing Failure Cases
Figure 7 illustrates two representative failure modes of UniFlow models in TruckScenes. In the first example, adverse weather introduces dense rain streaks that create spurious LiDAR returns, leading the model to produce localized artifacts. In the second example, extreme sparsity at long range causes the model not estimate flow for an object located nearly 100m away.